# The Normal Play of the Domination Game

### v3

@@ 3.3 Cycles
Figura 2 shows another example, where Alice wins by playing on the gray vertex in her first move, since this returns a position equivalent to the disjoint union of $P_{1}^{\prime}$, $P_{2}^{\prime}$, $P_{3}^{\prime}$, $P_{3}^{\prime\prime}$ and $P_{7}^{\prime\prime}$, with total nimber equal to $1\oplus 2\oplus 3\oplus 3\oplus 3=0$.
@@ 4 Partizan game in paths and cycles
In this section, we investigate for the first time the partizan variant of the Domination Game. As before, we focus on the normal variant, which we call Normal Partizan Domination Game. As done in the literature of combinatorial games, a partizan variant of an impartial game can be obtained by modifying the instances in such a way that the set of possible moves of Alice is disjoint from the set of possible moves of Bob. See for example the classical book Winning Ways for your Mathematical Plays from Berlekamp, Conway and Guy [1] and the games Hackenbush (impartial) and Blue-Red Hackenbush (partizan).
In the Normal Partizan Domination Game, the instance is a graph $G$ such that the vertices are colored either $A$ or $B$, and Alice (resp. Bob) can only select vertices colored $A$ (resp. $B$). As in the impartial variant (Normal Domination Game), the players alternately select vertices in such a way that every vertex selected in the game must dominate at least one vertex that was not dominated before. Moreover, the last to play wins in normal play.
@@
[…]
The Combinatorial Game Theory is a very rich theory with many results and definitions, such as infinitesimals, canonical form of a game, the Simplification Theorem for Numbers, Temperature Theory for hot games and so on. Here the main points are the following. A game is equal to 00 if and only if the first player loses, independently if the first player is Alice or Bob. Moreover, the sum $J_{1}+J_{2}$ of two games is the game resulting of the disjoint union of the games $J_{1}$ and $J_{2}$, that is, the player chooses either $J_{1}$ or $J_{2}$ and make a move there. Finally, every game $J=J_{1}+J_{2}$ has a unique canonical form $\{X|Y\}$, which can be determined in polynomial time given all options $J_{1}^{A}$ and $J_{2}^{A}$ of Alice and all options $J_{1}^{B}$ and $J_{2}^{B}$ of Bob.
The negative $-J$ of a game $J$ is the game in which the roles of Alice and Bob are changed. Notice that $J-J=J+(-J)=0$, since, for every move of the first player in one game, the opponent has the same move in the other game, which is then the last to play, winning the game.
@@ 4.2 Partizan game in paths
Let us apply the Combinatorial Game Theory on the Normal Partizan Domination Game in paths. We describe a path as its sequence of colors. For example, $AABA$ and $ABAA$ represents the same path $P_{4}$ $v_{1}v_{2}v_{3}v_{4}$ where $v_{1}$, $v_{2}$ and $v_{4}$ are colored A and $v_{3}$ is the only vertex colored B. Moreover, $BBAB=-AABA$, since the roles of Alice and Bob are changed. Notice that $AABA+BBAB=0$, that is, the first player loses in the disjoint union of two $P_{4}$’s with these colors.
As an example, let us determine the values of all paths with at most 4 vertices. For the instances of the path $P_{1}$, notice that $A=1$ and $B=-1$. For $P_{2}$, $AA=1$, $BB=-1$ and $AB=BA=\{0|0\}=*$.
For $P_{3}$, $AAA=2$ and, $BBB=-2$. Note that $AAB=BAA=\{\dot{A}B,0|A\dot{A}\}=\{*,0|1\}=\frac{1}{2}$, where $\dot{A}B$ means that the dotted vertex does not need to be dominated, but can be selected to dominate other vertices. Moreover, $ABB=BBA=-\frac{1}{2}$, $ABA=\{*|0\}=\downarrow$ and $BAA=\{0|*\}=\uparrow$.
Finally, for $P_{4}$, $AAAA=3$, $BBBB=-3$, $AAAB=BAAA=\{1|2\}=\frac{3}{2}$, $ABBB=BBBA=-\frac{3}{2}$, $AABA=ABAA=\left\{\frac{1}{2}|1\right\}=\frac{3}{4}$, $ABAA=BBAB=-\frac{3}{4}$ and $AABB=BBAA=ABAB=BABA=ABBA=BAAB=0$.
@@
If $i>j$, then $P_{i,j}=X_{i,j}=Y_{i,j}=Z_{i,j}=0$, since there is no vertex in these instances. If $i=j$, then $X_{i,j}=Y_{i,j}=Z_{i,j}=0$, since there is only one vertex in these instances and it is already dominated. If $i=j-1$, then $Z_{i,j}=0$, since there is only two vertices in this instance and both are already dominated.
@@
[…]
@@ 5 Conclusions and Open Problems
In the Normal Domination Game, it is possible to determine from Theorems 3.1 and 3.4 the nimber of any path $P_{n}$ and any cycle $C_{n}$, based only on the number $n$ (size of $\lceil\log_{2}n\rceil$ bits). Therefore, it is possible to determine the nimber of $G$ from the Sprague-Grundy theory by applying the bitwise-xor operation on the nimbers of the components of $G$, in logarithmic time. If the nimber is 00, the second player (Bob) wins. Otherwise, Alice wins.
@@
Other well known variant of combinatorial games is the misère variant, in which the last to play loses.
@@
The Misère Domination Game is PSPACE-complete in general graphs?
However, the Sprague-Grundy theory is a bit more complicated for misère games.

### v4

@@ 4.3 Cycles
Figura 3 shows another example, where Alice wins by playing on the gray vertex in her first move, since this returns a position equivalent to the disjoint union of $P_{1}^{\prime}$, $P_{2}^{\prime}$, $P_{3}^{\prime}$, $P_{3}^{\prime\prime}$ and $P_{7}^{\prime\prime}$, with total nimber equal to $1\oplus 2\oplus 3\oplus 3\oplus 3=0$.
5 Partizan variants are PSPACE-complete
In this section, we investigate for the first time the partizan variants of the Domination Game in normal play and misère play, which we call Normal Partizan Domination Game and Misère Partizan Domination Game. As done in the literature of combinatorial games, a partizan variant of an impartial game can be obtained by modifying the instances in such a way that the set of possible moves of Alice is disjoint from the set of possible moves of Bob. See for example the classical book Winning Ways for your Mathematical Plays from Berlekamp, Conway and Guy [1] and the games Hackenbush (impartial) and Blue-Red Hackenbush (partizan).
In the partizan variants of the domination game, the instance is a graph $G$ such that the vertices are colored either $A$ or $B$, and Alice (resp. Bob) can only select vertices colored $A$ (resp. $B$). As in the impartial variants, the players alternately select playable vertices. In normal (resp. misère) play, the last to play wins (resp. loses).
In this section, we prove the computational complexity of the partizan variants.
Theorem 5.1.
The Normal (resp. Misère) Partizan Domination Game is PSPACE-complete even in graphs with diameter two (resp. four).
Proof.
Consider an impartial domination game on normal play or misère play on a graph $H$ with at least one edge. Let $G$ be the graph obtained from $H$ by adding, for every vertex $v$ of $H$, a new vertex $v^{\prime}$ and the edges $vv^{\prime}$, $uv^{\prime}$ and $u^{\prime}v^{\prime}$ for every neighbor $u$ of $v$ in $H$. Color the vertices of $G$ in the following way: for every vertex $v$ of $H$, $v$ is colored $A$ and $v^{\prime}$ is colored $B$. Notice that, if $v$ (resp. $v^{\prime}$) is selected during the domination game, then $v^{\prime}$ (resp. $v$) is not playable anymore, since $v$ (resp. $v^{\prime}$) dominates the same vertices that can be dominated by $v^{\prime}$ (resp. $v$). This means that the normal and misère partizan domination games behave exactly like the impartial domination games. Therefore, since the diameter of $G$ is the same of $H$, then the Normal Partizan Domination Game is PSPACE-complete even in graphs with diameter two from Theorem 2.2 and the Misère Partizan Domination Game is PSPACE-complete even in graphs with diameter four from Theorem 3.1. ∎
@@ 6 Partizan game in paths and cycles
The games below are important and are denoted by 0, 11, 1-1, *, $\uparrow$ and $\downarrow$. In the empty game $0=\{|\}$, Alice and Bob have no moves and then the first to play loses. In the game $1=\{0|\}$, Bob has no move and any move of Alice wins, leading to a game 0. In the game $1=\{|0\}$, Alice has no move and any move of Bob wins, leading to a game 0. In the game star $\{0|0\}$, the first to play wins, leading to a game 0. Moreover, $\uparrow=\{0|*\}$ and $\downarrow=\{*|0\}$. The games $n$ and n-n can also be defined for every integer $n\geq 2$. In the game $n=\{n-1|\}$, Alice has $n$ more moves than Bob. In the game $n=\{|-(n-1)\}$, Bob has $n$ more moves than Alice.
The Combinatorial Game Theory is a very rich theory with many results and definitions, such as infinitesimals, canonical form of a game, the Simplification Theorem for Numbers, Temperature Theory for hot games and so on. Here the main points are the following. A game is equal to 0 if and only if the first player loses, independently if the first player is Alice or Bob. Moreover, the sum $J_{1}+J_{2}$ of two games is the game resulting of the disjoint union of the games $J_{1}$ and $J_{2}$, that is, the player chooses either $J_{1}$ or $J_{2}$ and make a move there. Finally, every game $J=J_{1}+J_{2}$ has a unique canonical form $\{X|Y\}$, which can be determined in polynomial time given all options $J_{1}^{A}$ and $J_{2}^{A}$ of Alice and all options $J_{1}^{B}$ and $J_{2}^{B}$ of Bob.
The negative J-J of a game $J$ is the game in which the roles of Alice and Bob are changed. Notice that JJ=J+(J)=0J-J=J+(-J)=0, since, for every move of the first player in one game, the opponent has the same move in the other game, which is then the last to play, winning the game.
@@ 6.2 Partizan game in paths
Let us apply the Combinatorial Game Theory on the Normal Partizan Domination Game in paths. We describe a path as its sequence of colors. For example, $AABA$ and $ABAA$ represents the same path $P_{4}$ $v_{1}v_{2}v_{3}v_{4}$ where $v_{1}$, $v_{2}$ and $v_{4}$ are colored A and $v_{3}$ is the only vertex colored B. Moreover, BBAB=AABABBAB=-AABA, since the roles of Alice and Bob are changed. Notice that $AABA+BBAB=0$, that is, the first player loses in the disjoint union of two $P_{4}$’s with these colors.
As an example, let us determine the values of all paths with at most 4 vertices. For the instances of the path $P_{1}$, notice that $A=1$ and B=1B=-1. For $P_{2}$, $AA=1$, BB=1BB=-1 and $AB=BA=\{0|0\}=*$.
For $P_{3}$, $AAA=2$ and, BBB=2BBB=-2. Note that $AAB=BAA=\{\dot{A}B,0|A\dot{A}\}=\{*,0|1\}=\frac{1}{2}$, where $\dot{A}B$ means that the dotted vertex does not need to be dominated, but can be selected to dominate other vertices. Moreover, $ABB=BBA=-\frac{1}{2}$, $ABA=\{*|0\}=\downarrow$ and $BAA=\{0|*\}=\uparrow$.
Finally, for $P_{4}$, $AAAA=3$, BBBB=3BBBB=-3, $AAAB=BAAA=\{1|2\}=\frac{3}{2}$, $BBB=BBBA=-\frac{3}{2}, AA$$ABA=ABAA=\left\{\frac{1}{2}|1\right\}=\frac{3}{4}$, $ABAA=BBAB=-\frac{3}{4}$ and $AABB=BBAA=ABAB=BABA=ABBA=BAAB=0$.
@@
If $i>j$, then $P_{i,j}=X_{i,j}=Y_{i,j}=Z_{i,j}=0$, since there is no vertex in these instances. If $i=j$, then $X_{i,j}=Y_{i,j}=Z_{i,j}=0$, since there is only one vertex in these instances and it is already dominated. If i=j1i=j-1, then $Z_{i,j}=0$, since there is only two vertices in this instance and both are already dominated.
@@
First let us solve for any path $P_{n}$. Algorithm 1 below is a dynamic programming algorithm that applies Lemma 6.1 in order to determine the value of $P_{1,n}$, which is the value of the entire instance. If $P_{1,n}>0$, Alice wins. If $P_{1,n}<0$, Bob wins. If $P_{1,n}=0$, the second player wins. If $P_{1,n}\ ||\ 0$, the first player wins.
@@ 7 Conclusions and Open Problems
In the Normal Domination Game, it is possible to determine from Theorems 4.1 and 4.4 the nimber of any path $P_{n}$ and any cycle $C_{n}$, based only on the number $n$ (size of $\lceil\log_{2}n\rceil$ bits). Therefore, it is possible to determine the nimber of $G$ from the Sprague-Grundy theory by applying the bitwise-xor operation on the nimbers of the components of $G$, in logarithmic time. If the nimber is 0, the second player (Bob) wins. Otherwise, Alice wins.
@@
Moreover, the Sprague-Grundy theory is a bit more complicated for misère games.
@@
Moreover, the Sprague-Grundy theory is a bit more complicated for misère games.
@@
The Misère Partizan Domination Game is polynomial time solvable in paths and cycles?